# The chanciness of time

### v1

@@ 1 Introduction
Digital networks for communications and computation pervade modern life, and scientists visualize networks to explain observable evidence, e.g. in sociology, biology, and physics. In networks, some events depend on others, and so occur temporally one after another, while other events are subject to no determined temporal order; one speaks of concurrent events. The mathematical expression that allows for concurrent events is a partial order. While, in theory, a mathematical partial order on a set can be extended to a total order [szpilrajn]; attempts to assign times to concurrent physical events encounter failures. For example, to determine that an event happened within a specific time interval, one has to decide that the event happened after the beginning of the time interval and before the end of it. The devices relied on for this decision—flip-flops—can teeter on edge. In the notorious case of asynchronous computer interrupts, we and others have measured the way that the instability of this teetering puts different parts of a digital network in conflict, with one part acting on a report that “$a$ is before bb” while another part acts on the report that “$a$ is after bb[76arbiter, 05aop, 19MST]. The network can then crash. More recently, we realized that determining temporal order in a three-way race depends on pairwise measurements, leading to a dramatic conclusion. In close calls the results for each pair are random. As a result, in a three-way race between concurrent events $e_{A}$, $e_{B}$, and $e_{C}$, a possible result is circular: $e_{A}<e_{B},e_{B}<e_{C}$, but $e_{C}<e_{A}$, in conflict with the transitivity of temporal ordering or even partial temporal ordering [19MST]. [^1] As a principal advisor to the National Institute of Science and Technology has put it, “time” as used in science and technology is an artifact [allan87, 14aop].
@@
To model concurrency in networks, we modify the mathematical definitions of token games played on marked graphs. Mathematical formulas used to represent motion do not themselves move; they sit still on the page; but they can represent snapshots of that motion. The current mathematics of marked graphs invokes “snapshots,” termed markings. A marking expresses the location of tokens at a moment of a play of a token game. If the players moved one at a time, as in chess, a play of the game could be recorded as a sequence of snapshots taken between moves, in moments when the positions of pieces on a game board are still. A marking of a marked graph presupposes such a snapshot, but for token games expressing a digital network, the moves are not “one at a time,” and there need be no moment for a snapshot to show a
@@
In section 2 we introduce history graphs along with their specializations appropriate to representing plays of token games. In section 3 we discuss token games played on game graphs to represent computational networks and their unpredictable changes, and we show a token game in which there can be no global snapshots. In section 4 we show how history graphs of plays of token games allow the study of these games to be based on local snapshots, without resort to global snapshots. In basing the mathematics of token games on local records, we highlight the acts of imagination necessary to interpret local snapshots as representing motion, acts reminiscent of those needed to infer the behavior of animals from tracks left in the sand.
@@ 2 Local records of interdependent events
We imagine local records from which to assemble the dependencies among events. An event depends directly on other events that we call its source events. Suppose that the record for each event names that event and also its source events, as in fig. 1(a). Each such record has a graphical representation, as in fig. 1(b) By tracing back sources of sources, the local records define a larger history graph, illustrated in fig. 1(c).
Figure 1: (a) Local records of event sources; (b) graphical portrayal of local records; (c) history graph assembled from local records.
@@
History graphs involve choosing a level of description. For example, a coarsening of history graph in fig. 1(c) can combine events $e_{6}$, $e_{7}$, and $e_{8}$ into the single event $e_{\displaystyle*}$. A coarsening of a history graph is expressed by a mapping of events and edges that breaks no edges and creates no cycles.
@@
Later on, we will consider records that include not just the names of sources, but also what they supply. For example, a source of an event of calculation can supply a label consisting of a string of bits (0’s and 1’s). Multiple history graphs can pertain to a given situation; for example, a smaller earlier history graph might be a subgraph of a larger history graph that expresses additional local records.
@@ 2.1 History graphs for networks of nodes
History graphs can represent locally recorded communication activity in a network. Then each event of the history graph represents one of the sequence of actions of a node of the network, so that the events of the history graph are partitioned into sets, one set of each node of the network. We speak of partitioned history graphs. We assume a finite set of nodes of the network, and we picture events of partitioned history graphs as arranged on columns, one column for each node of the network. Figure 3 shows an example.
@@
For an event in a partitioned history graph, we speak not only of source events and target events, but less specifically of source nodes and target nodes:
@@
The source node of an event in a partitioned history graph is the column containing its source event, and the target node of an event in a partitioned history is the column in which its target event lies.
In a partitioned history graph for a network undergoing changes during its operation, two events in the same column can differ in their source and target nodes.
@@ 2.2 Base graph of a partitioned history graph
For any partitioned history graph there is an associated base graph, defined as follows. The base graph has a node for each column of events of the history graph and an arrow between nodes if and only if the partitioned history graph has a corresponding edge. We call a mapping from a subgraph of a partitioned history graph to its base a projection if and only if:
@@ 3 Token games
A token game straddles mathematics and physical behavior. It is played by one or more players on a game board, according to rules, and yet the rules are sufficiently clear to be represented mathematically. The game board implements an image of a game graph, with a node for each player and arrows that express communication channels from node to node. [^4] As discussed in appendix A, the game graphs can have multiple arrows and loops. As illustrated in fig. 4, unlike a history graph, a game graph can contain cycles.
@@
For a token game to represent the logic of computations, two restrictions are necessary; the game must be live and safe. Liveness means that deadlocks are precluded.
@@ 3.1 Token game without any still moment for a global snapshot
Figure 5 illustrates a play of token game with unlabeled tokens, satisfying the firing rule, in which each node has still moments for local snapshots, but there is no “still moment” for all the nodes at once for a global snapshot to show unambiguous positions of tokens. A moment for a local snapshot at a node comes when the node is fireable but has not yet fired. The absence of any global “still moment” that motivates us to define marked graphs without invoking the notion of a marking, a.k.a. a global snapshot. The six pictures can be followed clockwise around and around, illustrating continuous play with no moment when all tokens are at rest. Hex(3,3), discussed more below, is one of three token games that can be played on the same “Hex” game graph.
@@ 3.2 Local records of three token games on the Hex game graph
A node can fire more than once, so the event name includes not only the node that fires but also an index that is incremented each time the node fires. The form of a local record is then [node name, index|| node name, index; node name, index;…], where events to the right of the “||” are source events. For example, table 1 shows local records for plays of three distinct token games that can be played on the game graph shown in fig. 4. The nodes of the graph are $v_{0},v_{1},v_{2},v_{3},v_{4}$ and $v_{5}$. The first record of the play of the game denoted by “Hex(1,5) is $v_{0}\mathpunct{:}1|v_{5}\mathpunct{:}0;v1\mathpunct{:}0$] . That record gives the sources of the event $v_{0}\mathpunct{:}1$ as $v_{5}\mathpunct{:}0$ and $v1\mathpunct{:}0$. The numbers after the colons are indices. The order in the list of records for each token game is arbitrary; in general the node firings are partially ordered, as determined by the indices. Note that the indices of any node indicate only the total order of firings of that node, so that any integer added to all the indices for a node has no significance.
@@ 4.1 History graphs for plays of live and safe token games
From one firing to the next firing of any node in a live and safe token game, the indices on local records for that node and its source nodes all increase by 1.
@@
For any node $v$ of the game graph with node $w$ as a source node, consider local records of any two node-firing events $v\mathpunct{:}m$ and $v\mathpunct{:}n$. Then the local records of firing events for $v$ show $v\mathpunct{:}m|w\mathpunct{:}m^{\prime};$…] and $v\mathpunct{:}n|w\mathpunct{:}n^{\prime};$…]. We argue that $n-m=n^{\prime}-m^{\prime}$. If not, there must be a different number of firings of $v$ than of $w$ connecting the two records of $v$ firings. But such a difference would require a $w$-firing that was not a source for a $v$-firing or a $v$-firing without a $w$-firing as a source event, both of which are precluded in a live and safe token game by the firing rule. Because consecutive firings of a node increment its firing index by 1, the indices of its source-node firings are also incremented by 1. ∎
@@ 4.1.1 Example
If a live token game is safe, and its history graph shows two successive firing events of a node $v$ of the game graph, there are directed paths from the first $v$-firing event through firing events of each of its targets and on to the second $v$-firing event.
@@
For a token game to be safe, the node $v$, having fired, cannot fire a second time before tokens have been removed from all its arrows for which it is the source node. This implies that the target nodes of $v$ must fire, removing tokens from the outgoing arrows of $v$, before that $v$ can fire again. Thus if $e_{v}$ is a firing event of $v$ and $e^{\prime}_{v}$ is the next firing event of $v$, each target event of $e_{v}$ is contained is in a path from $e_{v}$ to $e_{v}^{\prime}$. ∎
@@
To represent a computation by a live and safe token game, one must have a strong game graph, which precludes source and sink nodes. If a computation is presented as having a start and an end, suggesting source and sink nodes, we reframe the computation as in process-control: i.e., we augment the game graph with an “ENV” node for an environment, along with an arrow from each end node to this ENV node, and with an arrow from the ENV node to each start node [SRL].
@@ 5 History graphs to show changes in live and safe token games
Networks are often in flux, undergoing planned as well as unpredicted changes. We wondered how to represent the insertion of a new link into a chain of communications over a network without interrupting the chain? An answer was inspired by a study in biology that showed a series of snapshots as an amoeba inserted itself into a chain of amoeba forming a filament of a cellular slime mold. In [25bmb] we translated the series of snapshots into a series of live and safe marked graphs. Each next graph was related to the previous graph by a change such that the play of tokens could proceed without interruption [25bmb]. By means of the history graph illustrated in fig. 10, the node-firing events of a play that proceeds through such changes, predictable or not, can be portrayed not by a series of marked graphs, with their markings, but by a single history graph without invoking markings.
@@
In fig. 10, the numbers identify nodes, some of which are created and others of which are terminated during the changes that take place. In fig. 10(b) the creation and termination of nodes is visible along the verticle green dashed lines. The earliest node firings are for a 2-cycle (1,2) of node fiings representing an amoeba disjoint from the filament, which is characterize by the marked graph prior to step 1. This marked graph is characterized here by the nodes 3, 4, and 7. After the last change, step 5, the lone amoeba has inserted itself to from the graph to form a chain on the game graph characterized by the nodes 3, 5, 6, and 7. Details of the steps from game graph to game graph that preserve liveness and safety can be found in [25bmb].
@@
Here are the details. When $v_{3}\mathpunct{:}2$ became fireable, instead of firing it was separated into two nodes $v_{3}a$ and $v_{3}b$, each with one arrow pointing in and 1 arrow pointing out. Then $v_{3}a$ fires and is a source for $v2\mathpunct{:}3$ while $v_{3}b$ never fires (and so has no firing event in the history graph). The result of this splitting of the node and the firing of one part is to change the play from the Hex(3,3) game to the Hex(2,4) game. After the firing $v_{3}a\mathpunct{:}0$ the two parts of the $v_{3}$ node are rejoined. Note that the indices of any node indicate only the total order of firings of that node, so that any integer added to all the indices for a node has no significance.
@@ Acknowledgment
We thank Judith Clapp for incisive help in organizing our writing.
@@ Appendix A Appendix: Definitions
We say an arrow (for a game graph) or an edge (for a history graph) connects its source to its target. An arrow or an edge pointing into its target is called an in-arrow or in-edge, respectively; an arrow or an edge pointing out of its source is called an out-arrow or an out-edge.

### v2

@@ 1 Introduction
Digital networks for communications and computation pervade modern life, and scientists visualize networks to explain observable evidence, e.g. in sociology, biology, and physics. In networks, some events depend on others, and so occur temporally one after another, while other events are subject to no determined temporal order; one speaks of concurrent events. The mathematical expression that allows for concurrent events is a partial order. A partial order on a set of events orders an event with respect to some subset of events but not necessarily with respect to all the other events. For example, as shown in the figure (b) on the right, A, and B are unordered with respect to C, D, and E.
While, in theory, any mathematical partial order can be extended to a total order [szpilrajn]; attempts to assign times to concurrent physical events encounter failures. A striking example of trouble in assigning times has never been completely eliminated from digital computers, even after decades of experience. A computer takes steps, timed by a clock that ticks (faster than once per nanosecond). The computer has a memory, made of tiny flip-flops, into each of which a 0 or 1 can be written, doing away with whatever bit that was there before. The computer receives messages from other computers, coded into 0’s and 1’s. When a computer receives a 0 or 1 (written on an electronic token), it has to record whether the token came before or after a certain tick of its clock. So, what happens if the token carrying a 1 arrives just as the clock is ticking and the flip-flop holds a 0? Does the flip-flop change to 1 or stay with 0? The simple answer “It’s random.” turns out to be misleading. The flip-flop acts as a kind of hinge that could flip one way or flop the other way, so the hinge can hang up in the middle, teetering on edge, as illustrated below.
Teetering between ”0” and ”1”.
In the notorious case of asynchronous computer interrupts, we and others have measured the way that the instability of this teetering puts different parts of a digital network in conflict, with one part acting on a report that “$a$ is before bb” while another part acts on the report that “$a$ is after bb[76arbiter, 05aop, 19MST]. The network can then crash. More recently, we realized that determining temporal order in a three-way race depends on pairwise measurements, leading to a dramatic conclusion. In close calls the result for each pair is random. As a result, in a three-way race between concurrent events $e_{A}$, $e_{B}$, and $e_{C}$, a possible result is circular: $e_{A}<e_{B},e_{B}<e_{C}$, but $e_{C}<e_{A}$, in conflict with the transitivity of temporal ordering or even partial temporal ordering [19MST]. [^1] As a principal advisor to the National Institute of Science and Technology has put it, “time” as used in science and technology is an artifact [allan87, 14aop].
@@
To model concurrency in networks, we modify the mathematical definitions of token games played on marked graphs. Mathematical formulas used to represent motion do not themselves move; they sit still on the page; how something still can represent something moving seems mysterious until it strikes us that formulas represent snapshots of that motion. The current mathematics of marked graphs invokes “snapshots,” termed markings. A marking expresses the location of tokens at a moment of a play of a token game. If the players moved one at a time, as in chess, a play of the game could be recorded as a sequence of snapshots taken between moves, in moments when the positions of pieces on a game board are still. A marking of a marked graph presupposes such a snapshot, but for token games expressing a digital network, the moves are not “one at a time,” and there need be no moment for a snapshot to show a
@@
In section 2 we introduce history graphs along with their specializations appropriate to representing plays of token games. In section 3 we discuss token games played on game graphs to represent computational networks and their unpredictable changes, and we show a token game in which there can be no global snapshots. In section 4 we show how history graphs of plays of token games allow the study of these games to be based on local snapshots, without resort to global snapshots. In basing the mathematics of token games on local records, we highlight the acts of visualization necessary to interpret local snapshots as representing motion, acts reminiscent of those needed to infer the behavior of animals from tracks left in the sand.
@@ 2 Local records of interdependent events
We imagine local records from which to assemble the dependencies among events. A record of an event includes information that comes to it from other events. An event depends directly on other events that we call its predecessor events: If event $e$ is a predecessor of event $e^{\prime}$, we draw a directed edge from $e$ to $e^{\prime}$ ($e\rightarrow e^{\prime}$); also we speak of $e^{\prime}$ as a successor of $e$. Suppose that the record for each event names that event and also its predecessor events, as in fig. 1(a). Each such record has a graphical representation, as in fig. 1(b) By tracing back predecessors of predecessors, the local records define a larger history graph, illustrated in fig. 1(c).
Figure 1: (a) Local records of event predecessors; (b) graphical portrayal of local records; (c) history graph assembled from local records.
@@
History graphs involve choosing a level of granularity of description. For example, a coarsening of history graph in fig. 1(c) can combine events $e_{6}$, $e_{7}$, and $e_{8}$ into the single event $e_{\displaystyle*}$. A coarsening of a history graph is expressed by a mapping of events and edges that breaks no edges and creates no cycles.
@@
Later on, we will consider records that include not just the names of predecessors, but also what they supply. For example, a predecessor of an event of calculation can supply a label consisting of a string of bits (0’s and 1’s). Multiple history graphs can pertain to a given situation; for example, a smaller earlier history graph might be a subgraph of a larger history graph that expresses additional local records.
@@ 2.1 History graphs for networks of nodes
History graphs can represent locally recorded communication activity in a network. Then each event of the history graph represents one of the sequence of actions of a node of the network, so that the events of the history graph are partitioned into sets, one set of each node of the network. We speak of partitioned history graphs. We assume a finite set of nodes of the network. In our pictures of history graphs, we show events, drawn as small black squares, arranged in columns, indicated by vertical dashed lines in green, one column for each node of the network. Figure 3 shows an example.
@@
For an event in a partitioned history graph, we speak not only of predecessor events and successor events, but less specifically of predecessor nodes and successor nodes:
@@
A predecessor node of an event in a partitioned history graph is the column containing the corresponding predecessor event, and the successor node of an event in a partitioned history is the column containing the corresponding successor event.
In a partitioned history graph for a network undergoing changes during its operation, two events in the same column can differ in their predecessor and successor nodes.
@@ 2.2 Base graph of a partitioned history graph
For any partitioned history graph there is an associated base graph, defined as follows. The base graph has a node for each column of events of the history graph and an arrow between nodes if and only if the partitioned history graph has a corresponding edge.
As discussed in appendix A of the Appendix, we say an arrow of a game graph connects a source node to a target node, and an edge of a history graph connects a source event to a target event. We call a mapping from a subgraph of a partitioned history graph to its base a projection if and only if:
@@ 3 Token games
A token game straddles mathematics and physical behavior. It is played by one or more players on a game board, according to rules, and yet the rules are sufficiently clear to be represented mathematically. The game board implements a visual image of a game graph, with a node for each player and arrows that express communication channels from node to node. [^4] As discussed in appendix A, the game graphs can have multiple arrows and loops. As illustrated in fig. 4, unlike a history graph, a game graph can contain
cycles.
@@
For a token game to represent the logic of computations, two restrictions are necessary; the game must be live and safe.
@@ 3.1 Token game without any still moment for a global snapshot
Figure 5 illustrates a play of a token game with unlabeled tokens, satisfying the firing rule, in which each node has still moments for local snapshots, but there is no “still moment” for all the nodes at once for a global snapshot to show unambiguous positions of tokens. The six pictures represent global photographs attempting to catch markings. The small dotted circles are added to the photographs to show places on arrows where no tokens reside but which will eventually receive tokens from a firing that is in progress when the picture is taken. However, when those arrows receive tokens, other tokens will have been removed from other firings that have begun but are not yet complete. The six pictures can be followed clockwise around and around, illustrating continuous play with no moment when all tokens are at rest. It is the absence of any global “still moment” that motivates us to define marked graphs without invoking the notion of markings as global snapshots.
@@ 3.2 Local records of three token games on the Hex game graph
A node can fire more than once, so the event name includes not only the node that fires but also an index that is incremented each time the node fires. The form of a local record is then [node name, index|| node name, index; node name, index;…], where events to the right of the “||” are predecessor events. For example, table 1 shows local records for plays of three distinct token games that can be played on the game graph shown in fig. 4. The nodes of the graph are $v_{0},v_{1},v_{2},v_{3},v_{4}$ and $v_{5}$. The first record of the play of the game denoted by “Hex(1,5) is $v_{0}\mathpunct{:}1|v_{5}\mathpunct{:}0;v1\mathpunct{:}0$] . That record gives the predecessors of the event $v_{0}\mathpunct{:}1$ as $v_{5}\mathpunct{:}0$ and $v1\mathpunct{:}0$. The numbers after the colons are indices. The order in the list of records for each token game is arbitrary; in general the node firings are partially ordered, as determined by the indices. Note that the indices of any node indicate only the total order of firings of that node, so that any integer added to all the indices for a node has no significance.
@@ 4.1 History graphs for plays of live and safe token games
From one firing to the next firing of any node in a live and safe token game, the indices on local records for that node and its predecessor nodes all increase by 1.
@@
For any node $v$ of the game graph with node $w$ as a predecessor node, consider local records of any two node-firing events $v\mathpunct{:}m$ and $v\mathpunct{:}n$. Then the local records of firing events for $v$ show $v\mathpunct{:}m|w\mathpunct{:}m^{\prime};$…] and $v\mathpunct{:}n|w\mathpunct{:}n^{\prime};$…]. We argue that $n-m=n^{\prime}-m^{\prime}$. If not, there must be a different number of firings of $v$ than of $w$ connecting the two records of $v$ firings. But such a difference would require a $w$-firing that was not a predecessor for a $v$-firing or a $v$-firing without a $w$-firing as a predecessor event, both of which are precluded in a live and safe token game by the firing rule. Because consecutive firings of a node increment its firing index by 1, the indices of its predecessor-node firings are also incremented by 1. ∎
@@ 4.1.1 Example
If a live token game is safe, and its history graph shows two successive firing events of a node $v$ of the game graph, there are directed paths from the first $v$-firing event through a firing event of each of its successor events and on to the second $v$-firing event.
@@
For a token game to be safe, the node $v$, having fired, cannot fire a second time before tokens have been removed from all its arrows for which it is the source node. This implies that the successor nodes of $v$ must fire, removing tokens from the outgoing arrows of $v$, before that $v$ can fire again. Thus if $e_{v}$ is a firing event of $v$ and $e^{\prime}_{v}$ is the next firing event of $v$, each successor event of $e_{v}$ is contained is in a path from $e_{v}$ to $e_{v}^{\prime}$. ∎
@@
To represent a computation by a live and safe token game, one must have a strong game graph; a strong game graph allows no source (a node with not arrow pointing into it) and no sink (a node that lacks an arrow pointing out from it). If a computation is presented as having a start and an end, suggesting source(s) and sink(s), we reframe the computation as in process-control: i.e., we augment the game graph with an “ENV” node for an environment, along with an arrow from each end node to this ENV node, and with an arrow from the ENV node to each start node [SRL].
@@ 5 History graphs to show changes in live and safe token games
Networks are often in flux, undergoing planned as well as unpredicted changes. We wondered how to represent the insertion of a new link into a chain of communications over a network without interrupting the chain? An answer was inspired by a study in biology that showed a series of snapshots as an amoeba inserted itself into a chain of amoeba forming a filament of a cellular slime mold. In [25bmb] we translated the series of snapshots into a series of live and safe marked graphs. Each next graph was related to the previous graph by a change such that the play of the token game could proceed without interruption [25bmb]. By means of the history graph illustrated in fig. 10, the node-firing events of a play that proceeds through such changes, predictable or not, can be portrayed not by a series of marked graphs, with their markings, but by a single history graph without invoking markings.
@@
In fig. 10, the numbers identify nodes, some of which are created and others of which are terminated during the changes that take place. In fig. 10(b) the creation and termination of nodes is visible along the vertical green dashed lines. The earliest node firings are for a 2-cycle (1,2) of node firings representing an amoeba disjoint from the filament, which is characterize by the marked graph prior to step 1. This marked graph is characterized here by the nodes 3, 4, and 7. After the last change, step 5, the lone amoeba has inserted itself to from the graph to form a chain on the game graph characterized by the nodes 3, 5, 6, and 7. Details of the steps from game graph to game graph that preserve liveness and safety can be found in [25bmb].
@@
Here are the details. When $v_{3}\mathpunct{:}2$ became fireable, instead of firing it was separated into two nodes $v_{3}a$ and $v_{3}b$, each with one arrow pointing in and 1 arrow pointing out. Then $v_{3}a$ fires and is a predecessor for $v2\mathpunct{:}3$ while $v_{3}b$ never fires (and so has no firing event in the history graph). The result of this splitting of the node and the firing of one part is to change the play from the Hex(3,3) game to the Hex(2,4) game. After the firing $v_{3}a\mathpunct{:}0$ the two parts of the $v_{3}$ node are rejoined. Recall that the indices of any node indicate only the total order of firings of that node, so that any integer added to all the indices for a node has no significance.
@@ Acknowledgment
We thank Judith Clapp for incisive help in organizing our writing. We thank Susan Wood for helpful criticisms and suggestions.
@@ Appendix A Appendix: Definitions
An arrow (for a game graph) or an edge (for a history graph) connects its source to its target.